OpenStudy (anonymous):

(cos theta - sin theta)^2 -1=

OpenStudy (jdoe0001):

expand \(\bf [cos(\theta)-sin(\theta)]^2-1\) using the binomial theorem, see what you get

OpenStudy (anonymous):

multiply out and use the pythagorean identity: \(\sin^2\theta +\cos^2\theta = 1\)

OpenStudy (anonymous):

that whole thing is in parentheses ( ) which means sin - cos times sin - cos, or (sin theta - cos theta) squared, and I do not know what binomial theorem is.

OpenStudy (anonymous):

foil, (or distribute)

OpenStudy (anonymous):

thanks, but that was the wrong answer

OpenStudy (anonymous):

what did you get?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=1pSD8cYyqUo